40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16900     Entry time: Wed Jun 8 16:03:12 2022
Author: Koji 
Type: Update 
Category: VAC 
Subject: Power Outage 220608: Vacuum Recovery 

=== Observation ===

  • TP1/2/3 failed. AUX RP did not survive the power outage (of course)

=== Initial Recovery of TP2/TP3 ===

  • Stopped TP1/2/3
  • Restored RP AUX at the wall breaker module
  • Restarted TP2/3

=== Rough pumping down ===

  • Found RP1/RP3 left connected (unusual)
  • Turned on RP1 and RP3
  • Opened V6 and V7 for roughing of the pump spool. Now 0.3 torr for P2/P3/PA* (annuli) - ready to open V4 and V5 to backup TP1 by TP2 and TP3. Waiting for TP2/3 to be accelerated

=== Towards main volume pumping ===

  • TP2 accelerated and opened V4 (This caused V6 to close). RP1 and RP3 stopped.
  • TP3 hasitated to spin up. Stopped for now.
  • TP1 needed to be reset to clear the displacement errors. Once the errors were cleared, TP1 was restarted (acclerating).
  • TP1 at normal speed. Opened V1 (main volume pumping started)
     
  • Disconnected RP1/RP3 line.
     
  • TP3 was restarted and is now stable at 55k RPM (indicated as soft spin / low speed). Pushed "Stand By On" and "then "Stand By OFF". This makes the TP3 back to "NORMAL". Switched to "Stand By On".
  • TP2 was also switched to "Stand By On".

Now the main volume and the annuli are pumped down TP1 and TP2/TP3 with RP AUX backing.

The attachment is the pressure glitch for the main volume.

 

Attachment 1: Screenshot_2022-06-08_16-43-13.png  68 kB  Uploaded Wed Jun 8 17:43:29 2022  | Hide | Hide all
Screenshot_2022-06-08_16-43-13.png
ELOG V3.1.3-